OpenStudy (anonymous):

what 2 fractions are equivalent to -1/5

OpenStudy (anonymous):

what you think?

OpenStudy (anonymous):

here are a lot of them -1/5 = -2/10 = -3/15 = -4/20 = -5/25 = -6/30 = -7/35 = -8/40 = -9/45 = -10/50 = -11/55 = -12/60 = -13/65 = -14/70 = -15/75 = -16/80 = -17/85 = -18/90 = -19/95 = -20/100 how did i get them: If you want to find equivalent fractions to, say, 3/4, you can multiply the numerator and denominator by the same number. You don't change the value of a fraction if you multiply its top and bottom numbers by the same number because you are just multiplying by 1.

OpenStudy (anonymous):

@newwar248

OpenStudy (anonymous):

-2/10 -5/25

OpenStudy (anonymous):

i could be any one of them unless you have to pick.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

what are your choices?